# EtCon: Edit-then-Consolidate for Reliable Knowledge Editing

### v1

@@ 4.2 Knowledge Consolidation via Group Relative Policy Optimization
The reward function evaluates multiple aspects of the generated trajectory:
@@
This consolidation step effectively integrates the parametric knowledge acquired through TPSFT into the model’s CoT-based inference policy, ensuring that the edited facts are not merely memorized but can be coherently utilized in complex reasoning tasks while maintaining locality on unrelated inputs.
@@ 5.1 Experiment Settings
Datasets and Models. This work utilizes 1000 samples from each of three benchmark datasets, ZsRE Levy et al. (2017), COUNTERFACT Meng et al. (2022a), and QAEdit Yang et al. (2025b), to comprehensively evaluate the performance on knowledge editing tasks. We select two widely used LLMs, Llama-3-8B-Instruct Dubey et al. (2024) and Qwen-2.5-7B-Instruct Li et al. (2025a), as the base models for editing. For general ability evaluation, we use C-Eval Huang et al. (2023), CoQA Reddy et al. (2019), DROP Dua et al. (2019), SQuAD 2.0 Rajpurkar et al. (2018) and LogiQA Liu et al. (2020).
Baselines. We compare our method against two main categories: Parametric In-Place Editing methods (FT-M Zhang et al. (2024a), MEMIT Meng et al. (2022b), ALPHAEDIT Fang et al. (2024), MMKE Fu et al. (2025)) and External-Assisted Editing methods (WISE Wang et al. (2024a)). Parametric In-Place Editing methods are the main focus of this work, and we select the most representative methods in this category as baselines. For External-Assisted Editing methods, we select WISE as it is the SOTA method in this category.
Implementation Details. We conduct experiments using EasyEdit Xu et al. (2025) for evaluating various baselines, and employ the lm-evaluation-harness for assessing general capabilities. TPSFT is implemented through PSFT Zhu et al. (2025) for edit stage, while GRPO is built upon the EasyR1 Zheng et al. (2025b) for the consolidation stage. The specific hyperparameters are shown in Appendix A.1.
Evaluation Metrics. We evaluate our method along two principal axes: editing performance and general capability preservation. To assess editing performance, we employ the LLM-as-judge framework from Yang et al. (2025b); Gao et al. (2024); Gu et al. (2024b); Wang et al. (2024b), which mitigates the overestimation issue inherent in token-based metrics. In this framework, we leverage GPT-4.1 for a binary (correct/incorrect) evaluation of the model’s edited outputs to measure three key aspects: Reliability (edit success), Generalization (effectiveness on related inputs), and Locality (impact on unrelated inputs). To ensure that the editing process does not compromise the model’s broader abilities, we further evaluate its general capability preservation. To this end, we report Accuracy on the classification benchmarks C-Eval and LogiQA, alongside Exact Match (EM) and F1 scores for the question-answering datasets CoQA, DROP, and SQuAD 2.0. Details of real-world evaluation is in Appendix A.7
@@ 5.2 Main Results
Table 2 presents our evaluation of EtCon against existing baselines across three benchmarks under real-world lifelong editing evaluations. EtCon consistently outperforms all baselines across both model architectures. On Qwen-2.5-7B-Instruct, EtCon achieves 69.4% Reliability on ZsRE and 75.1% on QAEdit, surpassing the strongest baseline ALPHAEDIT by 53.5 and 75.1 percentage points respectively. Similar improvements occur on Llama-3-8B-Instruct, where EtCon reaches 73.5% Reliability on ZsRE versus FT-M’s 16.6%. Notably, EtCon maintains strong Generalization scores (60.8% on ZsRE, 63.0% on QAEdit for Qwen-2.5) while preserving acceptable Locality (24.2%-33.6%), confirming that our approach successfully preserves unrelated knowledge while performing targeted edits.
The local editing methods (MEMIT and ALPHAEDIT) fail catastrophically in lifelong editing. MEMIT collapses entirely on Qwen-2.5-7B-Instruct with near-zero performance across all metrics. ALPHAEDIT performs marginally better but remains highly unstable: it achieves 15.9% Reliability on ZsRE but completely fails on COUNTERFACT and QAEdit (0.0% across all metrics) for Qwen-2.5. Even when ALPHAEDIT reaches 61.0% Reliability on COUNTERFACT for Llama-3, its Locality drops to 16.1%, indicating severe knowledge disruption. This failure stems from destructive interference between sequential edits, where uncontrolled accumulation of weight deltas causes exponential growth in layer norms, leading to model collapse.
FT-M and WISE show improved stability over local editing methods but remain far below EtCon’s performance. FT-M achieves only 5.6% Reliability on ZsRE for Qwen-2.5 compared to EtCon’s 69.4%, while WISE performs even worse at 4.5%. On Llama-3, FT-M’s best result (27.9% on COUNTERFACT) still falls 39.2 percentage points below EtCon. These substantial performance gaps validate the effectiveness of our approach, which we attribute to two key design choices: the local editing in the edit stage preserves unrelated knowledge, and more crucially, the consolidation stage enables the reasoning network to effectively utilize the edited knowledge, thereby completing the critical final step in the knowledge editing pipeline. Extended experiments with up to 3,000 sequential edits demonstrating EtCon’s robustness in larger-scale lifelong editing scenarios are provided in Appendix A.3.
| ZsRE | COUNTERFACT | QAEdit | ||||||||||
| Method | Reli. | Gen. | Loc. | Reli. | Gen. | Loc. | Reli. | Gen. | Loc. | |||
Pre-edit | 4.4 | 3.2 | 28.5 | 1.0 | 0.5 | 36.9 | 9.8 | 10.1 | 36.2 | ||
| FT-M | 5.6 | 5.5 | 23.1 | 3.2 | 3.1 | 24.4 | 14.6 | 14.5 | 30.7 | |||
| MEMIT | 0.0 | 0.1 | 0.0 | 0.0 | 0.2 | 0.1 | 0.4 | 0.3 | 0.2 | |||
| ALPHAEDIT | 15.9 | 11.5 | 6.8 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | |||
| WISE | 4.5 | 3.3 | 19.1 | 1.4 | 1.5 | 31.0 | 7.1 | 9.7 | 16.9 | |||
| EtCon | 69.4 | 60.8 | 24.4 | 59.6 | 43.2 | 29.7 | 75.1 | 63.0 | 32.3 | |||
Pre-edit | 2.8 | 2.4 | 38.6 | 0.6 | 0.8 | 31.8 | 12.7 | 12.5 | 44.3 | ||
| FT-M | 16.6 | 15.5 | 29.3 | 27.9 | 18.6 | 10.5 | 34.1 | 33.2 | 30.1 | |||
| MEMIT | 0.1 | 0.1 | 0.0 | 0.3 | 0.7 | 0.4 | 0.2 | 0.7 | 0.0 | |||
| ALPHAEDIT | 18.7 | 14.0 | 6.3 | 61.0 | 43.8 | 16.1 | 18.2 | 14.9 | 7.5 | |||
| WISE | 4.3 | 3.1 | 2.2 | 1.3 | 0.8 | 31.3 | 8.1 | 13.3 | 0.9 | |||
| EtCon | 73.5 | 63.1 | 30.2 | 67.1 | 53.4 | 24.2 | 70.7 | 62.7 | 33.6 | |||
5.3 Analysis of Consolidation Stage
| DataSet | Metric | Base | FT-M |
MMKE |
ALPHA |
EtCon | ||||||
| Edited Knowledge | |||||||||||||||
| QAEdit | Reli. | 12.6 | 14.6 | 42.3 | 12.2 | 37.2 | 0.0 | 0.0 | 75.1 | ||||||
| Gen. | 13.9 | 14.5 | 34.1 | 10.4 | 31.4 | 0.0 | 0.0 | 63.0 | |||||||
| Loc. | 36.2 | 30.7 | 31.9 | 34.2 | 31.0 | 0.0 | 0.0 | 32.3 | |||||||
| General Capabilities | |||||||||||||||
| C-Eval | Acc. | 79.49 | 75.93 | 76.97 | 79.27 | 78.83 | 23.02 | 23.03 | 78.45 | ||||||
| CoQA | EM | 54.47 | 21.33 | 26.22 | 60.30 | 59.07 | 0.00 | 0.00 | 55.13 | ||||||
| F1 | 70.13 | 38.74 | 46.64 | 74.60 | 73.33 | 0.00 | 0.00 | 69.41 | |||||||
| DROP | EM | 2.21 | 2.37 | 2.79 | 10.30 | 8.46 | 0.00 | 0.00 | 2.52 | ||||||
| F1 | 9.94 | 13.31 | 14.59 | 24.32 | 21.59 | 0.00 | 0.00 | 8.60 | |||||||
| SQuAD | EM | 9.88 | 2.88 | 4.37 | 13.79 | 12.05 | 50.07 | 50.07 | 9.85 | ||||||
| F1 | 18.88 | 11.17 | 13.53 | 21.10 | 19.55 | 50.07 | 50.07 | 19.59 | |||||||
| LogiQA | Acc. | 38.71 | 37.02 | 37.79 | 41.01 | 39.17 | 21.81 | 21.81 | 38.40 | ||||||
To comprehensively evaluate the effectiveness of the proposed Consolidation phase, we conducted extensive experiments on the QAEdit dataset. We augmented three baseline knowledge editing methods (FT-M, MMKE, and ALPHAEDIT) with our Consolidation phase and compared their performance against our proposed EtCon method. As shown in Table 3, incorporating the Consolidation phase into FT-M and MMKE yields substantial improvements of 25-28% in Reliability and approximately 20% in Generalization metrics. These gains demonstrate that the Consolidation phase effectively bridges the gap between edited parametric knowledge and the model’s actual generation behavior, enabling successful knowledge utilization in real-world scenarios. Moreover, evaluations on multiple general-purpose benchmarks confirm that the Consolidation stage preserves the model’s general capabilities, with FT-M and MMKE maintaining their original performance levels and even exhibiting marginal improvements in certain cases. This preservation of general capabilities while enhancing editing performance validates the non-destructive nature of our consolidation mechanism.
However, the Consolidation phase cannot repair damage incurred during the editing stage. While FT-M with Consolidation achieves 5-8 percentage point improvements in EM and F1 scores on CoQA, these metrics remain substantially below the original model’s performance, highlighting the importance of careful knowledge updates during editing. MMKE’s design protects general capabilities but at the cost of reduced editing efficacy compared to EtCon. ALPHAEDIT exhibits model collapse after editing, which even the Consolidation phase cannot rectify. Overall, the Consolidation phase proves indispensable for knowledge editing, enabling effective generalization of newly edited knowledge while maintaining the model’s general capabilities. We further validate EtCon’s compatibility with reasoning-oriented architectures (DeepSeek-R1-Distill-Qwen-7B) in Appendix A.4. A detailed time efficiency analysis is provided in Appendix A.5.
5.4 ABLATION STUDIES
Stage Methods Reli. Gen. Loc. C-Eval CoQA sQuAD 2.0 Base - 0.6 0.8 31.8 50.82 78.20 29.52 Edit. w/ SFT 1.4 0.3 30.7 48.66 75.76 26.52 w/ TPSFT 3.3 1.8 30.2 50.07 78.52 34.60 Consolidate. w/o 56.1 22.4 24.7 - - - w/o 51.6 27.2 25.1 - - - Complete 67.1 53.4 24.2 - - -
We conduct a thorough ablation study on the COUNTERFACT dataset using Llama-3-8B-Instruct to isolate the individual contributions of each key component in our EtCon framework, with results presented in Table 4. In the Edit Stage, we compare our TPSFT against standard SFT. The results indicate that neither method alone is sufficient to enable reliable application of new knowledge, as reflected by low success and generalization scores. However, TPSFT demonstrates a clear advantage in preserving the model’s general capabilities, significantly mitigating the degradation observed with SFT. In the Consolidation Stage, building upon the TPSFT edit, we ablate components of our comprehensive reward function. Removing the cleanliness reward () causes a significant performance drop. Upon inspection, we find this encourages ”reward hacking,” where the model generates extraneous content to maximize its score, such as both old and new facts. The performance degrades more severely upon removing the consistency reward (), leading to catastrophic failures in reliability; for instance, the model might state the correct answer and then immediately contradict it. These findings confirm that our comprehensive reward function is critical for preventing such reward hacking and effectively steering the consolidation process toward reliable and coherent reasoning. For illustrative case studies, please refer to Appendix A.8.
@@ 6 Conclusion
In this paper, we identified that the critical gap between theoretical performance and practical effectiveness of knowledge editing methods stems from the absence of a consolidation stage that integrates parametric knowledge into the LLM’s actual generation behavior. To address this, we proposed the Edit-then-Consolidate (EtCon) framework, which combines Targeted Proximal Supervised Fine-Tuning (TPSFT) for precise knowledge editing with Group Relative Policy Optimization (GRPO) for effective consolidation. TPSFT updates targeted FFN weights within trust-region constraints to ensure reliable edits while preserving pretrained capabilities. Then GRPO aligns the edited knowledge with the model’s CoT-based inference policy through a comprehensive reward function. Our controlled experiments demonstrated the necessity of the knowledge consolidation stage, and comprehensive evaluations showed that EtCon significantly outperforms existing methods in lifelong editing scenarios, achieving efficient knowledge updates while maintaining locality and preserving model capabilities. These results suggest that explicitly decoupling editing and consolidation represents a promising paradigm for practical knowledge editing.

### v2

@@ 4.2 Knowledge Consolidation via GRPO
The reward function evaluates multiple aspects of the generated trajectory (as illustrated in Figure 4):
@@
This consolidation step integrates the parametric knowledge acquired through TPSFT into the model’s inference policy. By optimizing trajectory-level generation under autoregressive sampling rather than teacher-forcing, GRPO mitigates the train-inference mismatch, ensuring that edited knowledge is reliably elicited in actual behavior.
@@ 5.1 Experiment Settings
Datasets and Models. This work utilizes 1000 samples from each of four benchmark, ZsRE (Levy et al., 2017), COUNTERFACT (Meng et al., 2022a), MQuAKE-CF-v2 (Zhong et al., 2023) (1-edit subset), and QAEdit (Yang et al., 2025b), to evaluate the performance on knowledge editing tasks. We focus on the lifelong editing with sequential edits (batch size 1), which better reflects realistic deployment scenarios where knowledge updates arrive continuously over time. We select the Llama-3-8B-Instruct (Dubey et al., 2024) and Qwen2.5-7B-Instruct (Qwen et al., 2025), as the base models for editing. For general ability evaluation, we use C-Eval (Huang et al., 2023), CoQA (Reddy et al., 2019), DROP (Dua et al., 2019), SQuAD 2.0 (Rajpurkar et al., 2018) and LogiQA (Liu et al., 2020).
Baselines. We compare our method against two main categories: In-Place Parametric Edits methods (FT-M (Zhang et al., 2024a), MEMIT (Meng et al., 2022b), ALPHAEDIT (Fang et al., 2024), MMKE (Fu et al., 2025)) and External-Assisted Edits methods (WISE (Wang et al., 2024a), GRACE (Hartvigsen et al., 2023)). In-Place Parametric Edits methods are the main focus of this work, and we select the most representative methods as baselines.
Implementation Details. We conduct experiments using EasyEdit (Xu et al., 2025) for evaluating various baselines, and employ the lm-evaluation-harness (Gao et al., 2024a) for assessing general capabilities. TPSFT is implemented through PSFT (Zhu et al., 2025) for edit stage, while GRPO is built upon the EasyR1 (Zheng et al., 2025b) for the consolidation stage. More details are in Appendix A.1.
Evaluation Metrics. We evaluate our method along two principal axes: editing performance and pre-trained capability preservation. To assess editing performance, we employ the LLM-as-judge framework from (Yang et al., 2025b; Gao et al., 2024b; Gu et al., 2024b; Wang et al., 2024b), which mitigates the overestimation issue inherent in token-based metrics. Note that while GRPO training uses efficient rule-based rewards (§4.2), evaluation employs GPT-4.1 for semantic assessment to ensure rigor. Specifically, we conduct a binary (correct/incorrect) evaluation of the model’s edited outputs. For ZsRE, COUNTERFACT, and QAEdit, we measure Reliability (edit success), Generalization (effectiveness on related inputs), and Locality (impact on unrelated inputs). For MQuAKE-CF-v2, we report Edit-wise accuracy and Multi-hop accuracy. We report pretrained capability on C-Eval and LogiQA, and Exact Match (EM) and F1 on CoQA, DROP, and SQuAD 2.0. Real-world evaluation details are provided in Appendix A.6.
@@ 5.2 Main Results
Edit Efficiency. EtCon consistently achieves the best overall performance under real-world lifelong editing evaluation across benchmarks and base models. Table 1 presents comprehensive results on two LLMs. Across all six settings (3 benchmarks 2 base models), EtCon attains the top Avg. score, with large margins over the strongest competing method. On Qwen2.5-7B-Instruct, EtCon reaches Avg. of 51.5/44.2/56.8 on ZsRE/COUNTERFACT/QAEdit (improving the runner-up by +39.0/+30.9/+35.7). On Llama-3-8B-Instruct, EtCon further achieves Avg. of 55.6/48.2/55.7 (+35.1/+7.9/+23.2). These gains are primarily driven by substantially higher Reliability and Generalization, while maintaining competitive Locality (24.2%–33.6%), indicating that EtCon can internalize new facts and generalize them to paraphrases under sequential edits without overly compromising unrelated behaviors. We attribute this robustness to our two-stage design: TPSFT performs localized, stability-aware updates, and GRPO subsequently consolidates the edited knowledge under autoregressive trajectories.
| Editing Performance | General Capabilities | ||||||||||||
| Method | Reli. | Gen. | Loc. | Avg. | C-Eval | CoQA | DROP | SQuAD | LogiQA | Avg. | |||
| Acc. | EM | F1 | EM | F1 | EM | F1 | Acc. | ||||||
| Pre | 9.8 | 10.1 | 36.2 | 18.7 | 79.5 | 54.5 | 70.1 | 2.2 | 9.9 | 9.9 | 18.9 | 38.7 | 38.0 |
| FT-M | 14.6 | 14.5 | 30.7 | 19.9 | 75.9 | 21.3 | 38.7 | 2.4 | 13.3 | 2.9 | 11.2 | 37.0 | 25.3 |
| FT-M+Con. | 42.3 | 34.1 | 31.9 | 36.1 | 77.0 | 26.2 | 46.6 | 2.8 | 14.6 | 4.4 | 13.5 | 37.8 | 27.9 |
| MMKE | 12.2 | 10.4 | 34.2 | 18.9 | 79.3 | 60.3 | 74.6 | 10.3 | 24.3 | 13.8 | 21.1 | 41.0 | 40.6 |
| MMKE+Con. | 37.2 | 31.4 | 31.0 | 33.2 | 78.8 | 59.1 | 73.3 | 8.5 | 21.6 | 12.1 | 19.6 | 39.2 | 39.0 |
| ALPHA | 0.0 | 0.0 | 0.0 | 0.0 | 23.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.1 | 50.1 | 21.8 | 18.1 |
| ALPHA+Con. | 0.0 | 0.0 | 0.0 | 0.0 | 23.0 | 0.0 | 0.0 | 0.0 | 0.0 | 50.1 | 50.1 | 21.8 | 18.1 |
[…]
| Methods | Edit-wise | 2-hop | 3-hop | 4-hop |
| Base | - | 69.1 | 36.2 | 85.4 |
| FT-M | 19.1 | 2.7 | 3.8 | 4.2 |
| ALPHA | 20.4 | 13.9 | 11.3 | 9.5 |
| MMKE | 9.3 | 12.0 | 7.1 | 22.9 |
| MEMIT | 0.4 | 1.1 | 1.3 | 0.0 |
| EtCon | 54.1 | 24.2 | 14.6 | 33.3 |
| Stage | Methods | Reli. | Gen. | Loc. | GC-Avg |
| Base | - | 0.6 | 0.8 | 31.8 | 52.85 |
| Edit. | w/ SFT | 1.4 | 0.3 | 30.7 | 50.31 |
| w/ TPSFT | 3.3 | 1.8 | 30.2 | 54.40 | |
| Con. (after TPSFT) | w/o | 56.1 | 22.4 | 24.7 | 50.17 |
| w/o | 51.6 | 27.2 | 25.1 | 49.75 | |
| Complete | 67.1 | 53.4 | 24.2 | 50.25 |
| Edit Scope | Reli. | Gen. | Loc. | GC-Avg |
| Full Param. | 47.3 | 41.5 | 35.8 | 58.17 |
| Target edit | 75.1 | 63.0 | 32.3 | 55.83 |
Scaling to Longer Editing. Figure 5 shows performance evolution when scaling to 3,000 sequential edits on ZsRE. EtCon maintains stable performance: Reliability and Generalization decrease moderately from 73.5% to 58.2% and 63.1% to 48.7% respectively, while Locality remains within a narrow band (28%–32%). In contrast, FT-M degrades rapidly, with Reliability dropping from 16.6% to near-zero. These results demonstrate that EtCon remains stable even under longer editing sequences.
6 Analysis
6.1 Effectiveness of Consolidation Stage
Overall Performance of EtCon. As shown in Table 2, EtCon achieves the best trade-off between editing effectiveness and general capability preservation. EtCon attains 75.1% Reliability and 63.0% Generalization, substantially outperforming all baselines. Meanwhile, EtCon maintains general capabilities close to the original model (C-Eval: 78.5 vs 79.5, CoQA F1: 69.4 vs 70.1), demonstrating that our two-stage design successfully balances knowledge injection with pre-trained capability preservation.
Performance on Multi-hop Task. Table 4 further demonstrates EtCon’s advantage on MQuAKE-CF-v2. EtCon achieves 24.2% and 33.3% accuracy on 2-hop and 4-hop queries respectively, substantially outperforming FT-M (2.7%/4.2%). This indicates that our consolidation stage can propagate localized edits into more consistent multi-step inference. Under the EtCon paradigm, incorporating explicit reward for multi-hop queries during consolidation could further strengthen performance on multi-hop tasks.
Universality of Knowledge Consolidation. To evaluate the generality of our knowledge consolidation stage, we augment three baselines (FT-M, MMKE, and ALPHAEDIT) with Consolidation and compare them with EtCon on QAEdit. As shown in Table 2, adding Consolidation to FT-M and MMKE improves Reliability by 25% to 28% and Generalization by about 20%. These results suggest that Consolidation bridges the gap between edited parametric knowledge and autoregressive generation behavior, enabling more reliable knowledge utilization in practice.
Preservation of Pre-trained Capabilities. Evaluations across benchmarks show that the knowledge consolidation stage is non-destructive and preserves general capabilities. However, it cannot recover capabilities lost during editing: once ALPHAEDIT collapses, the knowledge consolidation stage cannot restore it. This indicates that reliable lifelong editing requires both lossless knowledge injection and post-edit consolidation. Future work under the EtCon paradigm should prioritize lossless knowledge editing.
6.2 Ablation Study
Edit Stage Ablation. We compare TPSFT with standard SFT in Table 5. The results show that neither method alone yields reliably editable behavior, as evidenced by consistently low Reliability and Generalization. Nevertheless, TPSFT offers a clear advantage in capability preservation, substantially reducing the degradation observed with SFT.
Consolidation Stage Ablation. We ablate key components of our consolidation reward (Table 5). Removing causes a clear drop (Reliability , Generalization ); inspection shows this encourages “reward hacking,” where the model emits extraneous content (e.g., both old and new facts) to inflate scores. Removing is even more detrimental to Reliability (), often producing catastrophic self-contradictions: the model states the correct answer and then immediately reverts. These ablations suggest that both reward components help mitigate reward hacking and improve output coherence; see Appendix A.8 for case studies.
Parameter Sensitivity. Figure 4 analyzes TPSFT hyper-parameters. Clipping threshold : Increasing improves Reliability and Generalization but reduces Locality, with a sharp drop at . Thus, provides the best trade-off, as larger allows stronger updates but increases policy drift. Reference update interval: Larger intervals consistently reduce Reliability and Generalization, while Locality remains nearly unchanged. Frequent updates (Interval) perform best because stale references increase the policy gap and over-activate clipping, limiting effective updates.
Effect of CoT Augmentation. Figure 6 compares TPSFT with and without CoT-augmented training labels. Without CoT augmentation, Reliability is slightly higher, but Locality drops sharply and general capabilities degrade (e.g., lower CoQA compared to base). This is because supervising only on the few target tokens leads to overfitting. CoT augmentation encourages the model to learn valid reasoning patterns rather than overfitting to isolated input-output mappings, thereby better preserving general capabilities.
Efficiency of Target Edit. We compare Target Edit, which updates a small set of targeted FFN weights, with Full Param update under the same TPSFT setup. As shown in Table 6, Target Edit yields substantially higher Reliability and Generalization, while incurring only a slight drop in general capability (GC-Avg). In contrast, Full Param attains notably weaker editing performance under the same optimization budget. We attribute this gap to reduced optimization efficiency under ratio clipping. With limited update steps per edit, Full Param spreads gradients across the whole model, weakening the learning signal on the edited fact. In addition, clipping may saturate early for many tokens, further shrinking effective updates. Overall, restricting TPSFT provides a more controllable and effective mechanism for knowledge editing.
@@ 7 Conclusion
In this work, we propose EtCon, an edit-then-consolidate framework for reliable knowledge editing. In Edit Stage, Targeted Proximal Supervised Fine-Tuning (TPSFT), performs constrained edits to inject knowledge while preserving pre-trained capabilities. In the Consolidate Stage, Group Relative Policy Optimization (GRPO), consolidates edits by aligning autoregressive generation with updated facts. Extensive experiments demonstrate that EtCon achieves superior editing reliability and generalization while preserving pre-trained capabilities in real-world editing scenarios.